OpenStudy (anonymous):

what's the answer for find the domain of f and f' for f(x)=(3-5x)^1/2

OpenStudy (sgadi):

let y=f(x) \[y^2=3-5x\] \[3-5x>0\] so domain for f is \[x\le 3/5\] I don't know for f'

OpenStudy (anonymous):

thanks for ur answer sgadi but what's d meaning of domain?

OpenStudy (sgadi):

domain is \[x\le\frac{3}{5}\]

OpenStudy (sgadi):

I have verified that for f' also the domain is same as f